Assume that the random variable X is normally distributed, with mean μ = 75 and standard deviation σ = 8 Compute the probability P(48 < X < 80).
Answer by Boreal(15235)

z=(x-mean)/sd
or (48-75)/8=-27/8 =-3.375
and (80-75)/8 or +0.625
probability is 0.7336
Can do this on calculator with 2nd VARS, Normalcdf(48,80,75,8)